Explain how you know that7/10 is greater than 1/2 but less than 4/5.
1
Expert's answer
2011-02-23T05:13:18-0500
1/2 *5/5& = 5/10
4/5 * 2/2 = 8/10.
Thus 5/10 < 7/10 and 8/10 > 7/10.